Juan scored five more 2-points shots than 3-points shots in his Friday basketball game. He scored two fewer free throw shots than 3-point shots. He scored 26 points in the game. How many of each shot did he make?
(This question totally has my stumped. I was thinking that it didn't have enough information, but then again, why would they have the question if one can't solve it.)
: Juan scored five more 2-points shots than 3-points shots in his Friday basketball game. He scored two fewer free throw shots than 3-point shots. He scored 26 points in the game. How many of each shot did he make?
(This question totally has my stumped. I was thinking that it didn't have enough information, but then again, why would they have the question if one can't solve it.)

Answer by checkley77(3654)

3X+2(X+5)+1(X-2)=26
3X+2X+10+X-2=26
6X=26-10+2
6X=18
X=18/6
X=3 3 POINT SHOTS WERE MADE.
3+5=8 2 POINT SHOTS WERE MADE.
3-2=1 FOUL SHOT WAS MADE.
PROOF:
3*3+2*8+1*1=26
9+16+1=26
26=26